# Towards universal property prediction in Cartesian space: TACE is all you need

### v1

@@ 2.6 Spectra — Water
Table 2: Relative RMSE (RRMSE, %) for dipole moment (D) and polarizability ($\alpha$) across different water systems. For each training set size, the best-performing results are highlighted in both bold and underline.00footnotetext: Results for all models except TACE are taken from refs. [13, 50]. For polarizability, some models predict the $\ell=0$ and $\ell=2$ components individually, and their errors are given as values separated by a slash. In this benchmark, we report errors on the validation set in accordance with previous studies.
@@
2.7 Hessians — GAP17 and GAP20
The GAP17 dataset [51] contains configurations representing a variety of carbon allotropes, including structures generated from AIMD simulations as well as those obtained via GAP-based active learning. The reference energies were computed using single-point DFT calculations within the LDA approximation, employing dense reciprocal-space meshes. The quantitative performance comparison is provided in Table 3, where our approach achieves superior accuracy with an energy RMSE of 8.55 meV/atom and forces RMSE of 299.88 meV/Å. Accurate modeling of vibrational properties has historically posed challenges for carbon systems in ref. [52]. Therefore, to further evaluate the curvature of the PES predicted by our model, we assess its capability to reproduce the phonon dispersion relations of diamond, following the protocol outlined in ref. [13]. Using the Phonopy Python package [53, 54], we computed the phonon spectrum of diamond and benchmarked our results against DFT calculations, as illustrated in Fig. 3. The phonon dispersion predicted by TACE demonstrates good agreement with DFT, capturing the essential vibrational characteristics. We noticed that HotPP [13] model trained on the GAP17 dataset, reported noticeable discrepancies in the high-frequency region near the $\Gamma$-point, which were attributed to limitations in the quality of the training data. In comparison, our model shows no significant deviations. To enable a more detailed comparison, we retrained our model using the more accurate and comprehensive GAP20 dataset [52], generated with the optB88-vdW exchange-correlation functional that includes dispersion interactions. The resulting phonon spectrum is in near-perfect agreement with DFT calculations. These results indicate that our model exhibits excellent extrapolation capability across both lower-level and higher-level DFT datasets.
@@ 2.7 Hessians — GAP17 and GAP20
2.8 Magnetic systems — FeAl
Magnetic materials play a key role in numerous technological applications. Compared with conventional 3N-dimensional PES, magnetic systems possess additional 3N degrees of freedom arising from magnetic moments, which pose greater demands on the modeling framework. Since 2020, models explicitly embedding magnetic moments have begun to emerge. The introduction of constrained DFT [57] in 2022 made the generation of datasets for magnetic systems feasible, and in 2024, MagNet [58] first utilized magnetic forces (i.e., the negative gradient of energy with respect to magnetic moments) as training labels. These advances now enable more thorough studies of magnetic systems. Leveraging the expressive power of our framework, we take the FeAl binary system as an example to demonstrate the performance of mTACE. The dataset used in this study was generated as described in ref.[59]. The errors on the validation set are summarized in Table 4. Beyond quantitative accuracy, mTACE is in principle capable of predicting formation energies and lattice parameters, reproducing thermal expansion behavior, and capturing total magnetic moments even at high Al concentrations, while ensuring robust structural and magnetic relaxations. In summary, mTACE offers a promising approach for investigating magnetic materials.
@@ 2.8 Magnetic systems — FeAl
mMTP [^1] Magnetic basis = 4.[59] mMTP [^2] Magnetic basis = 5.[59] mTACE
@@
2.9 External field — BaTiO3
The computational investigation of ferroelectric properties from first principles, particularly the dynamics of domain switching, is notoriously challenging due to the high computational cost. Here, we employ FITACE with the BaTiO3 dataset introduced in ref. [60], which consists of 59 frames containing 135 atoms each. The dataset was extracted from active-learning MD simulations in the limit of zero electric field and has been used to investigate the dynamic ferroelectric properties of BaTiO3 perovskite.
As a demonstration, we compare the errors obtained with our unified model against the official benchmark results (see Table 5). Our model achieves superior accuracy, which can be attributed both to the inherent advantages of the architecture and to its semi-local nature. This indicates that our framework is capable of faithfully reproducing various physical properties under external electric fields. We did not pursue further investigations, since models based on Cartesian or SO(3) spherical tensor products consistently show agreement between low error metrics and actual performance, making it unlikely for models to exhibit deceptively small errors while performing poorly in practice.
Table 5: MAE for electric enthalpy (U, meV/atom), forces (F, meV/Å), stress (S, $\mathrm{meV/\text{\AA }^{3}}$), polarization (P, $\mathrm{me\cdot\text{\AA }/atom}$), polarizability ($\alpha$, $\mathrm{me\cdot\text{\AA }^{2}\cdot V^{-1}}$) and Born effective charges (BEC, me) for BaTiO3. For models trained with 59 configurations, the best-performing results are highlighted in both bold and underline.22footnotetext: Results for Allegro-pol trained with 75 configurations are taken from the official paper [60], and it is unclear whether ZBL (Ziegler-Biersack-Littmark) potentials [61] were used. All other models were trained by us using the publicly available 59 configurations, while Allegro-pol was trained using the official input files. For the 59-configuration models, we explicitly indicate whether ZBL was applied.
@@ 2.10 Charged systems
The dataset constructed by Ko et al. [35] comprises structures in various charge states (neutral, positively, and negatively charged), where charge transfer, mediated by long-range electrostatic interactions, plays a significant role in determining the total energy. Notably, the dataset includes Ag trimers with total charge of +1 and -1, ionic Na9Cl8 clusters formed by the removal of a neutral Na atom, carbon chains terminated with hydrogen atoms in both neutral and positively charged states, and a periodic system consisting of a gold cluster adsorbed on a MgO(001) slab. Energy and forces were benchmarked to illustrate the effects of including charges and total charge embeddings and the charge errors correspond to models trained jointly on energy, forces, and charge labels. All models, except ACE [24] and CACE-LR [17], use LES or charge embeddings, either explicitly embedding charges or using charges as labels. We ensured charge conservation by distributing any excess atomic charge evenly across the system or using Qeq scheme. The errors reported in Table 6 correspond to the Qeq method. Overall, the results are highly satisfactory. One notable exception is the Ag systems, where our approach does not show a clear advantage in energy prediction. As long as embedding charges are used, incorporating the total charge as a continuous variable or omitting it has little effect on the error. One notable exception is the Ag systems, where our approach does not show a clear advantage in energy prediction. As long as embedding charges are used, incorporating the total charge as a continuous variable or omitting it has little effect on the error. We suspect this may be due to an inadequate embedding strategy for the total charge, a relatively low weight assigned to the energy term in the loss function, or possibly because competing models are overfitting, as suggested by the consistently low forces errors of our model.
Table 6: RMSE for energy (E, meV/atom), forces (F, meV/Å) and charges (Q, me) for different charged systems. The best-performing results are highlighted in both bold and underline.22footnotetext: Results for all models except TACE are taken from refs. [19, 35, 62].
@@
2.11 In-Domain — rMD17
The rMD17 dataset consists of ten small organic molecules whose configurations were initially generated by AIMD simulations in the MD17 dataset. To mitigate numerical noise, these configurations were re-computed at the PBE/def2-SVP level of theory, as detailed in ref. [64].
As shown in Table 7, mean absolute errors (MAE) results show that TACE not only improves the state of the art for several molecules but also demonstrates remarkable robustness across different molecular systems. Since multiple methods yield comparable accuracy on the standard task of predicting energy and forces using 950 configurations, we further evaluated TACE on a significantly reduced dataset containing only 50 structures. Even in this data-scarce scenario, TACE maintains high accuracy, highlighting its exceptional efficiency and strong generalization capability.
Table 7: MAE for energy (E, meV) and forces (F, meV/Å) for the rMD17 dataset. The best-performing results are both bolded and underlined.22footnotetext: Results for all models except TACE are from refs. [12, 65, 9].
$N_{\mathrm{train}}=950$ $N_{\mathrm{train}}=50$
System FCHL[65] Allegro[8] TensorNet[12] NequIP[7] MACE[9] TACE Nequip[7] MACE[9] TACE
aspirin E 6.2 2.18 2.4 2.3 2.2 1.88 19.5 17.0 14.86
F 20.9 7.3 8.9 8.2 6.6 6.05 52.0 43.9 41.99
azobenzene E 2.8 1.2 0.7 0.7 1.2 0.80 6.0 5.4 5.47
F 10.8 2.6 3.1 2.9 3 3.07 20.0 17.7 18.02
benzene E 0.3 0.3 0.02 0.04 0.4 0.15 0.6 0.7 0.29
F 2.6 0.2 0.3 0.3 0.3 0.35 2.9 2.7 1.89
ethanol E 0.9 0.4 0.5 0.4 0.4 0.39 8.7 6.7 6.82
F 6.2 2.1 3.5 2.8 2.1 2.46 40.2 32.6 32.43
malonaldehyde E 1.5 0.6 0.8 0.8 0.8 0.77 12.7 10.0 9.54
F 10.2 3.6 5.4 5.1 4.1 4.90 52.5 43.3 47.20
naphthalene E 1.2 0.2 0.2 0.9 0.5 0.30 2.1 2.1 2.02
F 6.5 0.9 1.6 1.3 1.6 1.35 10.0 9.2 9.02
paracetamol E 2.9 1.5 1.3 1.4 1.3 1.19 14.3 9.7 9.88
F 12.2 4.9 5.9 5.9 4.8 4.72 39.7 31.5 34.90
salicylic acid E 1.8 0.9 0.8 0.7 0.9 0.70 8.0 6.5 6.18
F 9.5 2.9 4.6 4 3.1 3.55 35.0 28.4 29.78
toluene E 1.6 0.4 0.3 0.3 0.5 0.33 3.3 3.1 2.68
F 8.8 1.8 1.7 1.6 1.5 1.51 15.1 12.1 12.98
uracil E 0.6 0.6 0.4 0.4 0.5 0.44 7.3 4.4 5.40
F 4.2 1.8 3.1 3.1 2.1 2.61 40.1 25.9 28.90
2.12 Out-of-Domain — 3BPA

### v2

@@ 2.5 Mixed Precision Training — Urea
Table 3: Relative RMSE, defined as the RMSE normalized by the intrinsic standard deviation of the validation set (RRMSE, %) for dipole moment (D) and polarizability ($\alpha$) across different water systems. For each training set size, the best-performing results are highlighted in both bold and underline.00footnotetext: Results for all models except TACE are taken from refs. [15, 45]. For polarizability, some models predict the $\ell=0$ and $\ell=2$ components individually, and their errors are given as values separated by a slash. In this benchmark, we report errors on the validation set in accordance with previous studies.
@@
2.6 Liquid water: structure, dynamics, and spectra
We assess the performance of TACE on liquid water using a dataset of 1593 configurations labeled at the revPBE0-D3 level  [46], which is known to provide a reliable description of the structure and dynamics of water across a range of pressures and temperatures [47]. Listed in Table 2, TACE attains the lowest RMSE of energy per atom and forces, comparing against invariant models (BPNN and DeepMD [3, 7]), linear ACE-style models, and recent equivariant message-passing models (NequIP, MACE and CAMP [8, 10, 17]). Notably, similar to ref. [43], a small-scale TACE model with invariant messages ($\mathrm{L_{max}=0}$) already outperforms top-performing models, indicating that the irreducible Cartesian framework is highly data-efficient and does not require large, high-rank tensor channels to reach state-of-the-art accuracy.
We then test whether this accuracy translates into physically correct dynamics and performed three independent 300-ps simulations of bulk water at 300 K, 500 K, 1000 K, 1500 K, and 2000 K using the $\mathrm{L_{max}=0}$ model. The O-O radial distribution function (RDF) is shown in Fig. 3 (a). At 300 K, the calculated RDF agrees remarkably well with experimental results [48]. We also computed the mean square displacement (MSD) and obtained the diffusion coefficient using the Einstein relation, yielding $0.26(1),1.58(1),3.17(7),4.66(7),5.90(28)\times 10^{-5}\ \mathrm{cm^{2}/s}$ at 300, 500, 1000, 1500, 2000K, which is in excellent consistency with the AIMD value of $2.67\times 10^{-5}\ \mathrm{cm^{2}/s}$ in 300K [49]. The TACE potential maintains stable MD simulations up to 2000 K, highlighting its robustness.
A central goal of this work is to explore whether TACE can simultaneously and accurately predict both scalar and tensorial properties. We therefore investigate whether TACE can describe dipole moments and polarizabilities of various water systems, and reproduce the infrared and Raman spectra of liquid water. For the dipole moment, we directly predict a vector $\nu=1$. For the polarizability, since it is a traceless and fully symmetric tensor, we predict two components: the trace part ($\ell=0$) and the traceless fully symmetric part ($\ell=2$). The prediction accuracy is reported in terms of the relative RMSE [44] in Table 3 and the explicit formula can be found in ref. [50]. TACE achieved very high accuracy in predicting dipole moments, achieving the best performance in three of the four systems considered, and shows competitive polarizability accuracy overall. However, some anomalous values can be observed in the Table 3, despite its strong performance across other systems. We believe that this is due to the fact that the polarizability labels are not true symmetric tensors; instead, they are constructed from the instantaneous dielectric tensor using the tensorial extension of Clausius-Mossotti relation [51]. HotPP fits it as a reducible Cartesian tensor by predicting a generic rank-2 tensor, which is later symmetrized, and includes an additional scalar corresponding to the diagonal element. By contrast, our model enforces full tensor symmetry by explicitly summing the $\ell=0$ and $\ell=2$ components. Thus, the slightly higher polarizability error is therefore best understood as a limitation of dataset labels [44] due to intrinsic noise, not of the representation itself.
To obtain the infrared and Raman spectra, we employed an NPT-equilibrated configuration of 512 water molecules from ref. [15]. This was followed by an additional 100 ps equilibration in the NVT ensemble using TACE trained on the dataset of ref. [52]. Subsequently, three production simulations were carried out in the NVE ensemble, each with a duration of 200 ps. Throughout all simulations, a time step of 0.5 fs was applied. The infrared and Raman spectra were then obtained by performing Fourier transforms of the respective autocorrelation functions. From Fig. 2, we can see that our computed results are in good agreement with experiments. The Raman spectrum shows a slight blue shift in the high-frequency region, which can also be attributed to the non-strict symmetry of the polarizability labels. At the same time, our spectra were obtained without applying broadening. It should be emphasized that quantum effects in water can significantly influence both the RDF and the simulated spectra, as demonstrated in refs. [53, 54, 55, 47]. Overall, our model achieves excellent performance in both stability and accuracy, while also demonstrating TACE’s effective simulation of scalars, tensors, and spectra.
@@ 2.6 Liquid water: structure, dynamics, and spectra
2.7 Hessians — GAP17 and GAP20
Accurate prediction of higher-order derivatives by the MLIP is crucial for phonon behavior, as the curvature of the true PES around equilibrium points directly determines the phonon frequencies and dispersion. We benchmarked on the GAP17 dataset [57] contains configurations representing a variety of carbon allotropes. On this dataset, TACE achieves an energy RMSE of 8.55 meV/atom and forces RMSE of 299.88 meV/Å, substantially outperforming other models in (Table 4). Accurate modeling of vibrational properties has historically posed challenges for carbon systems in ref. [58]. To directly assess curvature quality, we compute phonon dispersions for diamond from the TACE potential, following the same protocol used in ref. [15], and Phonopy [59, 60] to generate the phonon spectrum. As illustrated in Fig. 4, the phonon branches predicted by TACE agree well with DFT across the Brillouin zone. By contrast, we noticed that HotPP [15] model trained on the GAP17 dataset, reported noticeable discrepancies in the high-frequency region near the $\Gamma$-point. To enable a more detailed comparison, we retrained our model using the more accurate and comprehensive GAP20 dataset [58], generated with the optB88-vdW. The resulting phonon spectrum is in near-perfect agreement with DFT calculations, confirming that TACE can reproduce both forces and second-derivative information across lower-level and higher-level DFT datasets.
@@ 2.7 Hessians — GAP17 and GAP20
mMTP [^1] The results of mMTP correspond to magnetic basis = 4 and 5, respectively. [61] mMTP [^1] The results of mMTP correspond to magnetic basis = 4 and 5, respectively. [61] mTACE
@@
2.8 Magnetic systems — FeAl
Magnetic materials present a particular challenge for MLIPs because, in addition to the 3N atomic positional degrees of freedom that define a conventional potential energy surface, they also carry 3N spin degrees of freedom associated with local magnetic moments. MagNet [62] first utilized magnetic forces (i.e., the negative gradient of energy with respect to magnetic moments) as training labels. These advances now enable more thorough studies of magnetic systems.
Leveraging the universal embedding framwork, we extend TACE to mTACE, which treats per-atom magnetic moments as explicit vector-valued features and learns their interactions on the same footing as structural degrees of freedom. We evaluate mTACE on a FeAl binary system generated in ref.[61], as listed in Table 5. Beyond quantitative accuracy, mTACE is in principle capable of predicting formation energies and lattice parameters, reproducing thermal expansion behavior, and capturing total magnetic moments even at high Al concentrations, while ensuring robust structural and magnetic relaxations.
Table 6: MAE for electric enthalpy (U, meV/atom), forces (F, meV/Å), stress (S, $\mathrm{meV/\text{\AA }^{3}}$), polarization (P, $\mathrm{me\cdot\text{\AA }/atom}$), polarizability ($\alpha$, $\mathrm{me\cdot\text{\AA }^{2}\cdot V^{-1}}$) and Born effective charges (BEC, me) for BaTiO3. For models trained with 59 configurations, the best-performing results are highlighted in both bold and underline.00footnotetext: Results for Allegro-pol trained with 75 configurations are taken from the official paper [63], and it is unclear whether ZBL (Ziegler-Biersack-Littmark) potentials [22] were used. All other models were trained by us using the publicly available 59 configurations, while Allegro-pol was trained using the official input files. For the 59-configuration models, we explicitly indicate whether ZBL was applied.
@@ 2.8 Magnetic systems — FeAl
2.9 External field — BaTiO3
Modeling ferroelectric perovskites such as BaTiO3 under external electric fields is challenging for first-principles methods. To assess whether TACE can learn such field-coupled responses, we use its field-embedded variant (FITACE), which incorporates the external electric field as an equivariant through the universal embedding framwork.
We trained FITACE with the BaTiO3 dataset introduced in ref. [63], which consists of 59 frames extracted from MD simulations in the limit of zero electric field. We compare the unified model with the benchmark results (see Table 6), and FITACE achieves greater accuracy due to its architecture and semi-local nature. This indicates that our framework is capable of faithfully reproducing various physical properties under external electric fields. When we additionally include a short-range ZBL term (FITACE-ZBL), the accuracy further improves and remains competitive with, or better than, Allegro-pol trained on a larger set of 75 configurations. These results indicate that the universal embedding framework makes TACE extendable, with minimal modification, to systems under external electric fields.
@@ 2.10 Charged systems
We evaluate charge-aware learning on the dataset constructed by Ko et al. [38], which inculdes structures in various charge states where long-range electrostatics and charge transfer are decisive. Ag trimers with total charge of +1 and -1, ionic Na9Cl8 clusters formed by the removal of a neutral Na atom, carbon chains terminated with hydrogen atoms in both neutral and positively charged states, and a periodic system consisting of a gold cluster adsorbed on a MgO(001) slab.
Utilizing the universal embedding framework, energy and forces were benchmarked to illustrate the effects of including charges and total charge embeddings and the charge errors correspond to models trained jointly on energy, forces, and charge labels. All models, except ACE [23] and CACE-LR [19], use LES or charge embeddings, either explicitly embedding charges or using charges as labels. Overall, the results are highly satisfactory, shown in Table 7. One notable exception is the Ag systems, where our approach does not show a clear advantage in energy prediction. As long as embedding charges are used, incorporating the total charge as a continuous variable or omitting it has little effect on the error. We suspect this may be due to an inadequate embedding strategy for the total charge, a relatively low weight assigned to the energy term in the loss function, or possibly because competing models are overfitting, as suggested by the consistently low forces errors of our model.
Table 7: RMSE for energy (E, meV/atom), forces (F, meV/Å) and charges (Q, me) for different charged systems. The best-performing results are highlighted in both bold and underline.00footnotetext: Results for all models except TACE are taken from refs. [21, 38, 64].
@@
Table 8: Average distances of main chain carbon, carboxylic oxygen, and anhydride oxygen atoms from the Ag(111) surface (in Å).00footnotetext: The benchmark result of the TACE model was obtained with a batch size of 16 at the 216th epoch and MAE are compared against DFT.
uMLIPs finetune from-scratch
EXP DFT OrbV3 7Net MACE NequIP Allegro MACE NequIP TACE
[66] [67] con-inf MF-omat OMAT-medium OAM-L OAM-L OMAT-medium RECIO[68] LES
main chain carbon 2.86 2.97 3.38 3.22 3.21 3.73 3.51 3.00 2.98 2.97
carboxylic oxygen 2.66 2.67 3.07 3.03 3.40 3.29 2.79 2.72 2.68 2.69
anhydride oxygen 2.98 2.85 3.18 3.15 3.51 3.37 3.02 2.99 2.92 2.86
MAE - - 0.38 0.30 0.54 0.63 0.28 0.07 0.03 0.01
Table 9: Energy barriers (eV) for CO oxidation on various PdAg surfaces. Failures in searching TS are indicated by ×.00footnotetext: The benchmark result of the TACE model was obtained with a batch size of 16 at the 216th epoch and MAE are compared against DFT.
uMLIPs finetune from-scratch
System DFT OrbV3 MACE NequIP Allegro MACE NequIP TACE
[68] con-inf OMAT-medium OAM-L OAM-L OMAT-medium RECIO[68] LES
Pd100 0.63 0.60 0.66 0.65 0.88 0.69 0.67 0.65
Pd110 0.69 0.64 0.69 0.55 0.56 0.70 0.79 0.67
Pd111 1.39 × × × × × 1.35 1.29
Pd-TB 0.86 0.71 1.04 1.03 1.02 1.17 0.91 1.12
Pd-cube 1.28 1.01 1.17 1.17 1.35 1.23 1.18 1.23
Pd-S/Ag(111) 0.42 0.46 × × × × 0.42 0.38
Pd-line/Ag(111) 0.55 × × 0.64 × 0.51 0.58 0.51
Pd1/Ag(111) 0.28 0.27 × 0.40 0.31 × 0.38 0.29
Pd4/Ag(111) 0.23 × 0.28 × 0.32 × 0.21 0.20
Pd9/Ag(111) 0.87 × × × × × 0.73 0.87
Pd13/Ag(111) 0.74 × 0.36 × 0.60 0.45 0.60 0.54
Pd20/Ag(111) 0.90 0.69 0.89 0.85 0.76 0.92 0.99 0.91
MAE - 0.11 0.11 0.10 0.15 0.11 0.07 0.07
Table 10: Energy barriers (eV) for various Pd(111) and PdAg surfaces with different adsorbates. Failures in searching TS are indicated by ×.00footnotetext: The benchmark result of the TACE model was obtained with a batch size of 16 at the 216th epoch and MAE are compared against DFT.
uMLIPs finetune from-scratch
System DFT OrbV3 MACE NequIP Allegro MACE NequIP TACE
[68] con-inf OMAT OAM-L OAM-L OMAT RECIO[68] LES
Pd(111)-C2H3 0.87 0.72 0.74 0.74 0.55 0.78 0.73 0.79
Pd(111)-C2H4 0.79 0.70 0.76 0.68 0.68 0.72 0.74 0.71
Pd(111)-C2H5 0.79 0.67 0.61 0.63 × 0.74 0.73 0.85
Pd(111)-C2H6 0.71 × 0.50 0.44 0.54 0.53 0.76 0.70
1 ML-Pd(111)-C2H3 0.67 × 0.72 0.75 × × 0.74 0.36
1 ML-Pd(111)-C2H4 0.69 0.65 0.65 0.65 0.57 0.64 0.62 0.57
1 ML-Pd(111)-C2H5 0.90 × 0.60 0.67 0.59 0.72 0.64 0.75
1 ML-Pd(111)-C2H6 0.45 0.45 0.39 0.56 0.39 0.45 0.42 0.48
0.25 ML-PdAg3(111)-C2H3 0.58 × 0.61 0.62 × 0.60 0.51 0.63
0.25 ML-PdAg3(111)-C2H4 0.37 0.59 0.52 0.48 0.51 0.47 0.44 0.43
0.25 ML-PdAg3(111)-C2H5 0.42 0.67 0.61 0.66 0.62 0.68 0.71 0.62
0.25 ML-PdAg3(111)-C2H6 0.17 0.33 × 0.34 0.28 0.25 0.35 0.32
1 ML-PdAg(111)-C2H3 0.58 0.52 0.52 0.53 0.43 0.52 0.48 0.49
1 ML-PdAg(111)-C2H4 0.36 × 0.67 0.42 0.70 0.45 0.43 0.35
1 ML-PdAg(111)-C2H5 0.79 0.78 0.67 0.66 0.65 0.72 0.46 0.70
1 ML-PdAg(111)-C2H6 0.36 0.44 0.41 0.50 0.40 0.39 0.46 0.45
MAE - 0.11 0.13 0.13 0.17 0.09 0.12 0.10
2.11 Advantages and Limitations of ACE — REICO-AgPdCHO
Prior benchmarks use well-established databases designed to probe fundamental E/F/S potential energy surfaces or specific physical embeddings. These datasets most contains high-symmetry MD-derived structures which often clusters around near-equilibrium state, on which, as we shown, TACE achieves very high accuracy. Beside what we already shown, MLIP also also pursue generality, which is governed to a large extent by the chemical-space diversity of the training data. Recent efforts such as RECIO [68] sampling learns from randomized structures, reducing dependence on preselected configuration manifolds and pushing MLIPs toward DFT-like generality; likewise, recent new public datasets like OMat24[69], also emphasizes non-equilibrium coverage. Yet while such non-standard or non-system-specific data broaden coverage, they also raise the bar for training. On the AgPdCHO dataset [68], for example, ACE-based models often require extensive hyperparameter tuning; otherwise training can exhibit a spontaneous RMSE rise and eventually fail (even when customized metric scaling or Huber loss is applied). By contrast, NequIP trained reliably across most settings, suggesting an architecture-data interaction rather than a purely data issue. We hypothesize that the randomly generated structures are hard to be simultaneously captured well lack enough physical constraints, which leads to conflicts with the strong inductive biases of ACE-based frameworks. In practice, we also observe that two-layer ACE-based models can, in some cases, become more sensitive to outliers.
As a solution, we integrate LES as a plug-in long-range correction, which supplies appropriate physical constraints, that are especially helpful for models with short receptive fields, while remaining fully compatible with TACE’s design. With LES add-on, TACE can effectively learn from the complex datasets generated by the RECIO method, overcoming the inherent limitation of ACE framework. We benchmark the TACE trained AgPdCHO EMLP(element-based machine learning potentials) with selected test cases from the original work that are representative in heterogeneous catalysis, including adsorption of PTCDA on Ag(111) (Table 8.), CO oxidation on various PdAg surfaces (Table 9), and Acetylene hydrogenation on Pd(111), H covered Pd(111) and AgPd alloys (Table 10), Baker reaction (see SI Table S1). We employed the Atomic Simulation Environment package [70] to perform the following structure optimizations and transition state (TS) searches. We have also included the latest uMLIP (MACE-OMAT-medium [40], OrbV3-conservative-inf-omat [71], 7Net-omat [72], NequIP-OAM-L, Allegro-OAM-L [73]) and fine-tuned MACE-OMAT-medium with REICO data for comparison. Overall, we find TACE delivers the most balanced, zero-shot performance among all models, sub-0.03 Å deviations for PTCDA/Ag(111) adsorption distances, MAE $\approx$ 0.07 eV for CO-oxidation barriers across diverse Pd/PdAg with 100% TS success, and within $\leq$ 0.1 eV of DFT across coverage/alloy conditions. We verified the reliability of the predicted TS geometries by confirming their structural correctness and ensuring that each valid TS exhibited exactly one imaginary frequency (as the Hessians were available, frequencies could be obtained either by finite-difference calculations or by diagonalization after coordinate transformation, with the latter introducing less numerical noise). For OrbV3-conservative-inf-omat, the model failed to obey the appropriate physical constraints. As a result, the models may encounter issues such as energy non-conservation in the NVE ensemble [74], failure to locate transition state pathways [75], or even extremely unusual and completely incorrect optimization trajectories (e.g., adsorption optimization of PTCDA using OrbV3-direct-inf; see SI Fig. S2 ). It is also interesting to see the fine-tuned MACE-OMAT did not yield consistent gains and, in some cases, underperformed its pretrained counterpart. Our preliminary conclusion is that the conflict between ACE and RECIO persists, the randomized RECIO data contain a higher fraction of distributional outliers, which can confuse the model and degrade generalization.
2.12 uMLIP — TACE-MatPES
Here, we demonstrate the extrapolative capabilities of our model using a relatively small uMLIP dataset. In this case, we restrict the product basis with $\nu_{1}\leq\nu_{2}$” to further improve computational efficiency, and we train two types of models: an element-aware model and an element-agnostic model. Apart from the element-dependent part in the product basis, the rest of the architecture is kept identical. During the training of the element-aware model, we observed that its performance on the shear modulus test was more sensitive to the presence of outliers. After a certain number of epochs, although the overall accuracy continued to improve, abnormal values appeared for a few outlier samples. For fairness, we did not filter out these test results. Instead, we defined the best epoch as the point occurring within the first 20% of epochs before such outliers began to emerge. For consistency, we adopted the exact same data partitioning scheme as used in the MatPES publications [76] of CHGNet [77], M3GNet [78], and TensorNet [14]. Using this setup, we evaluate the performance of the TACE-MatPES and compare its results on more comprehensive benchmarks. It can be seen that, whether using a small-parameter model or an extremely large-parameter model, we achieve superior results. Notably, TACE offers great flexibility in controlling model parameters; for model covering 89 elements, it allows us to freely choose between 1M and 30M parameters, and even smaller or larger parameter counts are permissible. For further comparisons regarding model size, hyperparameters, and speed, please refer to SI 2.1.